Let T = f(x, y) be the temperature at the point (r, y) on the circle r = cost, y-inf.0≤ t≤2 and suppose that


ar 8-4y 8y-42


. Find where the maximum and minimum temperatures on the circle occur by examining the derivatives dT> dt and dT/dt


r=4x^ 2 -4xy+4y^ 2 . . Suppose that Find the maximum and minimum values of T7 on the circle


1
Expert's answer
2021-11-25T23:26:48-0500

"\frac{\partial }{\partial x}\left(T\right)=8x-4y\\
\frac{\partial }{\partial y}\left(T\right)=8y-4x\\
x = \cos t \\
y= \sin t\\
x^2+y^2=1\\
\sin ^2\left(x\right)+\cos ^2\left(x\right)=1\\
g = x^2 +y^2\\
d∇g = 2x i +2yj\\
=∇+= \lambda ∇ g \\
∇+= (8x-4y)i + (8y -4x)j \\
∇h= (8x-4y)i+(8y-4x)j + \lambda (2x i +2y j)\\
∇h=0\\
8 x-4y + 2x \lambda = 0....(1)\\
dy -4x + 2y \lambda = 0.....(2)\\
2x λ=4y-8x\\
λ= \frac{4y-8x}{2x}\\
2x λ+ 8x = 4y\\
x(2 λ+8)=4y\\
x= \frac{4y}{2λ+8}\\
Aslo\\
2y λ+8y = 4x\\
y(2λ+8) =4x\\
y= \frac{4x}{2λ+8}\\
\left(\frac{4y}{2λ+8}\right)^2+\left(\frac{4x}{2λ+8}\right)^2=1\\
\left(\frac{1}{2λ+8}\right)^2\left(16y^2+16x^2\right)=1\\
Thus \\
\left(\frac{4}{2λ+8}\right)^2=1\\
\frac{4}{\left(λ+4\right)^2}=1\\
4=\left(λ+4\right)^2\\
λ=-2,\:λ=-6\\
When \space λ= -2, x = y \\
When \space λ= -6, x = -y \\
\implies x = ± \frac{1}{\sqrt{2}}\\
Maxima \implies x = - \frac{1}{\sqrt{2}}, y = \frac{1}{\sqrt{2}}\\
Manima \implies x = \frac{1}{\sqrt{2}}, y = \frac{1}{\sqrt{2}}\\
Min_T= 4(\frac{1}{\sqrt{2}})^2-4(\frac{1}{\sqrt{2}})(\frac{1}{\sqrt{2}})+4(\frac{1}{\sqrt{2}})=2\\
Max_T= 4(-\frac{1}{\sqrt{2}})^2-4(\frac{1}{\sqrt{2}})(-\frac{1}{\sqrt{2}})+4(\frac{1}{\sqrt{2}})=6"